# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 4.1. Uncountability of almost disjoint families
Observe that $\dim(W)=M$, so if $U\subseteq W$ is a subspace such that $\dim(U)>\frac{M}{2}$, then $U$ intersects every proper subspace of $M$ non-trivially. Since $\{V_{k}:k<n\}$ is a collection of proper subspaces of $M$ with pairwise trivial intersections with at least $2$ elements, each $V_{k}$ is at most $\frac{M}{2}$-dimensional. Consequently, every $U_{i}$ contains at most $N^{\frac{M}{2}}$ vectors, which implies that:
$\left|\bigcup_{k<n}V_{k}\right|\leq n\cdot N^{\frac{M}{2}}.$
On the other hand, $|W|=N^{M}$. Therefore, we may choose $M$ large enough so that $N^{M}>n\cdot N^{\frac{M}{2}}$, contradicting the fact that $W=\bigcup_{k<n}U_{k}$.
@@
For the general case, suppose that $\{V_{k}:k<n\}$ is almost disjoint and $E=\bigcup_{k<n}V_{k}$. Let $M$ be large enough so that $M>\max(\operatorname{supp}[V_{j}\cap V_{k}])$ for all $j\neq k$. Let $W^{\prime}:=\operatorname{span}\{e_{n}:n\geq M\}$, and we have that $W^{\prime}=\bigcup_{k<n}W^{\prime}\cap V_{k}$. Then $\{W^{\prime}\cap V_{k}:k<n\}$ is a collection of infinite-dimensional subspaces of $W^{\prime}$ with pairwise trivial intersections, contradicting the argument above. ∎
@@
We show now construct some $V=\operatorname{span}\{x_{m}:m<\omega\}\subseteq E$ which intersects $V_{k}$ trivially for every $k<n$. Let $x_{0}\in E\setminus\bigcup_{k<n}V_{k}$ be any vector. Now suppose that we have defined $x_{i}$ for $i<m$ such that for all $k<n$:
$\operatorname{span}\{x_{i}:i<m\}\cap V_{k}=\{0\}.$
@@
For each $k<n$, let $U_{k}:=V_{k}+\operatorname{span}\{x_{i}:i<m\}$. By the claim in the proof of Lemma 4.4, $\{U_{k}:k<n\}$ is an almost disjoint family of subspaces of $E$, so by applying Claim 6 to $\{U_{k}:k<n\}$, $E\neq\bigcup_{k<n}U_{k}$. Let $x_{m}\in E\setminus\bigcup_{k<n}U_{k}$. Note that $\{x_{i}:i\leq m\}$ is linearly independent, as $x_{i}\in U_{0}$ for all $i<m$ but $x_{m}\notin U_{0}$. We shall show that:
$\operatorname{span}\{x_{i}:i\leq m\}\cap V_{k}=\{0\}.$
@@
If we can complete this induction step, then we are done as $V\cap V_{k}=\bigcup_{m<\omega}\operatorname{span}\{x_{i}:i<m\}\cap V_{k}=\{0\}$ for all $k<n$.
Let $y\in\operatorname{span}\{x_{i}:i\leq m\}\cap V_{k}$ for some $k<n$. Then $y=\sum_{i\leq m}\lambda_{i}x_{i}$ for some $\lambda_{i}\in\mathbb{F}$. If $\lambda_{m}\neq 0$, then we have that:
$x_{m}=\lambda_{m}^{-1}\left(y-\sum_{i<m}\lambda_{i}x_{i}\right)\in U_{k},$
contradicting the choice of $x_{m}$. Therefore, $\lambda_{m}=0$, which implies that:
$y\in\operatorname{span}\{x_{i}:i<m\}\cap V_{k}=\{0\},$
@@ 4.2. Ideals and semicoideals generated by almost disjoint families
If $\mathcal{A}$ is an almost disjoint family of $\omega]^{\omega}$, then the following families of sets are often studied:
@@
It is simple to verify that $\mathcal{I}^{+}(\mathcal{A})=\mathcal{I}^{++}(\mathcal{A})$ iff $\mathcal{A}$ is maximal. Motivated by this, Smythe introduced a version of $\mathcal{H}(\mathcal{A})$ for almost disjoint families of block subspaces. $\mathcal{H}(\mathcal{A})$ is a semicoideal, and Smythe studied the conditions that $\mathcal{A}$ must satisfy to make $\mathcal{H}(\mathcal{A})$ a full semicoideal.
@@
Definition 4.7.
Let $\mathcal{A}$ be an almost disjoint family of infinite-dimensional subspaces. We then define the following:
$\mathcal{I}(\mathcal{A})$ $:=\left\{Y\subseteq E:\begin{matrix}Y\setminus(\langle A_{0}\rangle\cup\cdots\cup\langle A_{n-1}\rangle)\text{ is small}\\ \text{for some }A_{0},\dots,A_{n-1}\in\mathcal{A}\end{matrix}\right\},$
$\mathcal{I}^{++}(\mathcal{A})$ $:=\{Y\subseteq E:\{A\in\mathcal{A}:Y\cap\langle A\rangle\text{ is big}\}\text{ is infinite}\},$
@@
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an infinite almost disjoint family, and let $Y\subseteq E$. The following are equivalent:
  1. (1)
    $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$.
  2. (2)
    $\langle Y\rangle\in\mathcal{I}^{++}(\mathcal{A})$ or there exists some $B\in E^{[\infty]}$, with $\langle B\rangle\subseteq\langle Y\rangle$, such that $B$ is almost disjoint from every element of $\mathcal{A}$.
Consequently, if $\mathcal{A}$ is mad, then $\mathcal{H}^{-}(\mathcal{A})=\mathcal{H}(\mathcal{A})$.
Proof.
We remark that for any infinite almost disjoint family $\mathcal{A}$, $\mathcal{I}^{++}(\mathcal{A})\subseteq\mathcal{I}^{+}(\mathcal{A})$.
[…]
We conclude this part with three useful lemmas we shall use repeatedly in this paper.
Lemma 4.9.
Let $A\in E^{[\infty]}$ and $Y\subseteq E$. Then $\langle A\rangle\cap\langle Y\rangle$ is small iff $\langle A/N\rangle\cap\langle Y\rangle=\emptyset$ for some $N$. Consequently:
@@
Suppose that $\langle A\rangle\cap\langle Y\rangle$ is big. Let $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle A\rangle\cap\langle Y\rangle$. For any $N$, there exists some $M$ such that $B/M\leq A/N$. This implies that $\langle B/M\rangle\subseteq\langle A/N\rangle\cap\langle Y\rangle$, so in particular, $\langle A/N\rangle\cap\langle Y\rangle\neq\emptyset$ for all $N$.
Conversely, suppose that $\langle A/N\rangle\cap\langle Y\rangle\neq\emptyset$ for all $N$. We construct some $B=(y_{n})_{n<\omega}\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle A\rangle\cap\langle Y\rangle$ inductively as follows: Assuming that $y_{0},\dots,y_{n-1}$ have been defined so far, let $N$ be large enough so that $\operatorname{supp}(y_{n-1})<A/N$. We let $y_{n}\in\langle A/N\rangle\cap\langle Y\rangle$ be any element. Then $\langle B\rangle\subseteq\langle A\rangle\cap\langle Y\rangle$, so $\langle A\rangle\cap\langle Y\rangle$ is big. ∎
@@
Let $\mathcal{A}\subseteq E^{[\infty]}$ be an almost disjoint family, and let $Y\subseteq E$. Suppose further that $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$ and there does not exist any $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle Y\rangle$ and $B$ is almost disjoint from every element of $\mathcal{A}$. Then there exists an infinite set $\{A_{n}:n<\omega\}\subseteq\mathcal{A}$ and an infinite disjoint family $\{B_{n}:n<\omega\}$ such that for all $n$:

### v5

@@ 4.1. Uncountability of almost disjoint families
Observe that $\dim(W)=M$, so if $U,U^{\prime}\subseteq W$ are two subspaces such that $\dim(U),\dim(U^{\prime})>\frac{M}{2}$, then $U$ and $U^{\prime}$ intersect non-trivially. Since $\{U_{k}:k<n\}$ is a collection of proper subspaces of $W$ with pairwise trivial intersections, and $n>2$, there is at most one $k$ such that $\dim(U_{k})>\frac{M}{2}$. We may assume that if such a $k$ exists, then $\dim(U_{0})>\frac{M}{2}$. Note that since $U_{k}$ is non-trivial for $k>0$ and $U_{0}\cap U_{k}\neq\emptyset$, we have that $\dim(U_{0})\leq M-1$. Consequently, for each $k>0$, $U_{k}$ contains at most $N^{\frac{M}{2}}$ vectors, which implies that:
$\left|\bigcup_{k<n}U_{k}\right|\leq(n-1)\cdot N^{\frac{M}{2}}+N^{M-1}.$
On the other hand, $|W|=N^{M}$. Since:
$N^{M}-((n-1)N^{\frac{M}{2}}+N^{M-1})$ $=(N-1)N^{M-1}-(n-1)N^{\frac{M}{2}}$
$\geq N^{M-1}-(n-1)N^{\frac{M}{2}}$
$=(N^{\frac{M}{2}-1}-n+1)N^{\frac{M}{2}}$
$>0\text{ for $M$ large enough}$
we have that $W\supsetneq\bigcup_{k<n}U_{k}$, a contradiction.
@@
For the general case, suppose that $\{V_{k}:k<n\}$ is almost disjoint and $E=\bigcup_{k<n}V_{k}$. Let $M$ be large enough so that $M>\max(\operatorname{supp}(x))$ for all $x\in V_{j}\cap V_{k}$ and $j\neq k$. Let $W^{\prime}:=\langle e_{n}:n\geq M\rangle$, and we have that $W^{\prime}=\bigcup_{k<n}W^{\prime}\cap V_{k}$. Then $\{W^{\prime}\cap V_{k}:k<n\}$ is a collection of infinite-dimensional subspaces of $W^{\prime}$ with pairwise trivial intersections, contradicting the argument above. ∎
@@
We shall now construct some $V=\langle x_{m}:m<\omega\rangle\subseteq E$ which intersects $V_{k}$ trivially for every $k<n$ by defining $x_{m}$ inductively. Let $x_{0}\in E\setminus\bigcup_{k<n}V_{k}$ be any vector. Now suppose that we have defined $x_{i}$ for $i<m$ such that for all $k<n$:
$\langle x_{i}:i<m\rangle\cap V_{k}=\{0\}.$
@@
For each $k<n$, let $U_{k}:=V_{k}+\langle x_{i}:i<m\rangle$. By the claim in the proof of Lemma 4.4, $\{U_{k}:k<n\}$ is an almost disjoint family of subspaces of $E$, so by applying Claim 6 to $\{U_{k}:k<n\}$, $E\neq\bigcup_{k<n}U_{k}$. Let $x_{m}\in E\setminus\bigcup_{k<n}U_{k}$. Note that $\{x_{i}:i\leq m\}$ is linearly independent, as $x_{i}\in U_{0}$ for all $i<m$ but $x_{m}\notin U_{0}$. We shall show that:
$\langle x_{i}:i\leq m\rangle\cap V_{k}=\{0\}.$
@@
If we can complete this induction step, then we are done as $V\cap V_{k}=\bigcup_{m<\omega}\langle x_{i}:i<m\rangle\cap V_{k}=\{0\}$ for all $k<n$.
Let $y\in\langle x_{i}:i\leq m\rangle\cap V_{k}$ for some $k<n$. Then $y=\sum_{i\leq m}\xi_{i}x_{i}$ for some $\xi_{i}\in\mathbb{F}$. If $\xi_{m}\neq 0$, then we have that:
$x_{m}=\xi_{m}^{-1}\left(y-\sum_{i<m}\xi_{i}x_{i}\right)\in U_{k},$
contradicting the choice of $x_{m}$. Therefore, $\xi_{m}=0$, which implies that:
$y\in\langle x_{i}:i<m\rangle\cap V_{k}=\{0\},$
@@ 4.2. Ideals and semicoideals generated by almost disjoint families
If $\mathcal{A}$ is an almost disjoint family in $\omega]^{\omega}$, then the following families of sets are often studied:
@@
It is simple to verify that $\mathcal{I}^{+}(\mathcal{A})=\mathcal{I}^{++}(\mathcal{A})$ iff $\mathcal{A}$ is maximal. Motivated by this, Smythe introduced a version of $\mathcal{H}(\mathcal{A})$ for almost disjoint families of subspaces. $\mathcal{H}(\mathcal{A})$ is a semicoideal, and Smythe studied the conditions that $\mathcal{A}$ must satisfy to make $\mathcal{H}(\mathcal{A})$ a full semicoideal.
@@
Note that if $V\subseteq E$ is a subspace, then $V$ is big iff $V$ is infinite-dimensional.
If $E$ is a vector space over $\mathbb{F}_{2}$, then by Hindman’s theorem, every small set is very small. On the other hand, if $E$ is a vector space over $\mathbb{F}\neq\mathbb{F}_{2}$, then this is no longer true — fix any $\mu\in\mathbb{F}\setminus\{0,1\}$. We may define the set:
$Y:=\{x\in E:x=e_{n}+y\text{ for some $e_{n}<y$ or $y=0$}\}.$
Observe that $Y$ and $Y^{c}$ are both small — for any $A\in E^{[\infty]}$, let $x\in\langle A\rangle$ be any non-zero vector. Then $x=\lambda_{n}(x)e_{n}+y$ for some $e_{n}<y$ or $y=0$ and $\lambda_{n}(x)\neq 0$, so $\frac{1}{\lambda_{n}(x)}x\in\langle A\rangle\cap Y$ and $\frac{\mu}{\lambda_{n}(x)}x\in\langle A\rangle\cap Y^{c}$. However, since $E=Y\cup Y^{c}$ and $E$ is big, $Y$ is not very small.
@@
Let $A\in E^{[\infty]}$ and let $V\subseteq E$ be a subspace. Then $\langle A\rangle\cap V$ is small iff $\langle A/N\rangle\cap V=\emptyset$ for some $N$. Consequently:
@@
Suppose that $\langle A\rangle\cap V$ is big. Let $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle A\rangle\cap V$. For any $N$, there exists some $M$ such that $B/M\leq A/N$. This implies that $\langle B/M\rangle\subseteq\langle A/N\rangle\cap V$, so in particular, $\langle A/N\rangle\cap V\neq\emptyset$ for all $N$.
Conversely, suppose that $\langle A/N\rangle\cap V\neq\emptyset$ for all $N$. We construct some $B=(y_{n})_{n<\omega}\in E^{[\infty]}$ such that $\langle B\rangle\subseteq\langle A\rangle\cap V$ inductively as follows: Assuming that $y_{0},\dots,y_{n-1}$ have been defined so far, let $N$ be large enough so that $\operatorname{supp}(y_{n-1})<A/N$. We let $y_{n}\in\langle A/N\rangle\cap V$ be any element. Since $\langle A\rangle\cap V$ is a subspace, we have that $\langle B\rangle\subseteq\langle A\rangle\cap V$, so $\langle A\rangle\cap V$ is big. ∎
Lemma 4.8.
Let $U,V\subseteq E$ be subspaces.
  1. (1)
    If $U\cap V$ is small, then $U\cap V$ is very small.
  2. (2)
    If $U\cap V$ is small, then $U\setminus V$ is very small.
Consequently, if $U$ is infinite-dimensional, then it is not possible to have both $U\cap V$ and $U\setminus V$ to be small.
Proof.
Assume that $U\cap V$ is small. Let $Y\subseteq E$, and suppose that $\langle A\rangle\subseteq(U\cap V)\cup Y$ for some $A=(x_{n})_{n<\omega}\in E^{[\infty]}$. Since $U\cap V$ is a finite-dimensional space, we may let $N$ be large enough so that $\operatorname{supp}(x_{N})>y$ for all $y\in U\cap V$. Then $\langle A/N\rangle\cap(U\cap V)=\emptyset$, so $\langle A/N\rangle\subseteq Y$. Therefore, $Y$ is big.
Now assume that $U\setminus V$ is small. Let $Y\subseteq E$, and suppose that $\langle A\rangle\subseteq(U\setminus V)\cup Y$. We consider three cases. For the first case, suppose that $\langle A\rangle\cap U$ is small. By Lemma 4.7, let $N$ be large enough so that $\langle A/N\rangle\cap U=\emptyset$. Then this implies that $\langle A/N\rangle\subseteq Y$, so $Y$ is big.
For the second case, suppose that $\langle A\rangle\cap V$ is big. Let $B\in E^{[\infty]}$ be such that $\langle B\rangle\subseteq\langle A\rangle\cap V$. Then $\langle B\rangle\subseteq(U\setminus V)\cup Y$, and since $\langle B\rangle\cap(U\setminus V)=\emptyset$, we have that $\langle B\rangle\subseteq Y$. Therefore, $Y$ is big.
For the last case, suppose that $\langle A\rangle\cap U$ is big but $\langle A\rangle\cap V$ is small. Let $B\in E^{[\infty]}$ be such that $\langle B\rangle\subseteq\langle A\rangle\cap U$, and note that $\langle B\rangle\cap V$ is also small. By Lemma 4.7, let $N$ be large enough so that $\langle B/N\rangle\cap V=\emptyset$. Since $\langle B/N\rangle\subseteq U$, this implies that $\langle B/N\rangle\subseteq U\setminus V$, contradicting our assumption that $U\setminus V$ is small. ∎
Definition 4.9.
Let $\mathcal{A}$ be an almost disjoint family of subspaces. We then define the following:
$\mathcal{I}(\mathcal{A})$ $:=\left\{Y\subseteq E:\begin{matrix}Y\setminus(V_{0}\cup\cdots\cup V_{n-1})\text{ is small}\\ \text{for some }V_{0},\dots,V_{n-1}\in\mathcal{A}\end{matrix}\right\},$
$\mathcal{I}^{++}(\mathcal{A})$ $:=\{Y\subseteq E:\{V\in\mathcal{A}:Y\cap V\text{ is big}\}\text{ is infinite}\},$
Lemma 4.10.
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces. Then $\mathcal{I}^{++}(\mathcal{A})\subseteq\mathcal{I}^{+}(\mathcal{A})$ (so $\mathcal{H}(\mathcal{A})\subseteq\mathcal{H}^{-}(\mathcal{A})$).
Proof.
Let $Y\in\mathcal{I}(\mathcal{A})$, and let $V_{0},\dots,V_{n-1}\in\mathcal{A}$ be such that $Y\setminus(V_{0}\cup\cdots\cup V_{n-1})$ is small. Then for any $U\in\mathcal{A}$ such that $U\neq V_{i}$ for any $i<n$, we have that:
$Y\cap U$ $\subseteq\left(Y\setminus\bigcup_{i<n}V_{i}\right)\cup\left(U\cap\bigcup_{i<n}V_{i}\right)$
$=\left(Y\setminus\bigcup_{i<n}V_{i}\right)\cup\bigcup_{i<n}V_{i}\cap U.$
By Lemma 4.8, $V_{i}\cap U$ is very small for all $i<n$, so $\bigcup_{i<n}V_{i}\cap U$ is very small. Since $Y\setminus\bigcup_{i<n}V_{i}$ is small, $Y\cap U$ is small. Therefore:
$\{U\in\mathcal{A}:Y\cap U\text{ is big}\}\subseteq\{V_{0},\dots,V_{n-1}\},$
so $Y\notin\mathcal{I}^{++}(\mathcal{A})$. ∎
Lemma 4.11.
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces, and let $U\subseteq E$ be a subspace. The following are equivalent:
  1. (1)
    $U\in\mathcal{I}^{+}(\mathcal{A})$.
  2. (2)
    $U\in\mathcal{I}^{++}(\mathcal{A})$ or there exists some $B\in E^{[\infty]}$, with $\langle B\rangle\subseteq U$, such that $B$ is almost disjoint from every element of $\mathcal{A}$.
Consequently, if $\mathcal{A}$ is mad, then $\mathcal{I}^{+}(\mathcal{A})=\mathcal{I}^{++}(\mathcal{A})$ (hence $\mathcal{H}^{-}(\mathcal{A})=\mathcal{H}(\mathcal{A})$).
Proof.
[…]
We conclude this part with three useful lemmas that we shall use repeatedly in this paper.
@@
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces, and let $U\subseteq E$ be a subspace. Suppose further that $U\in\mathcal{I}^{+}(\mathcal{A})$ and there does not exist any $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq U$ and $B$ is almost disjoint from every element of $\mathcal{A}$. Then there exists an infinite set $\{V_{n}:n<\omega\}\subseteq\mathcal{A}$ and an infinite disjoint family $\{B_{n}:n<\omega\}$ such that for all $n$, $\langle B_{n}\rangle\subseteq V_{n}\cap U$. .
Proof.
[…]
Lemma 4.14.
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces, and let $U\subseteq E$ be a subspace. If $U\in\mathcal{I}^{+}(\mathcal{A})$, then there exists some $B\in\mathcal{H}^{-}(\mathcal{A})$ such that $\langle B\rangle\subseteq U$. If furthermore $U\in\mathcal{I}^{++}(\mathcal{A})$, then $B$ may be chosen such that $B\in\mathcal{H}(\mathcal{A})$.
Proof.
If $U\notin\mathcal{I}^{++}(\mathcal{A})$, then by Lemma 4.11 there exists some $B\in E^{[\infty]}$ such that $\langle B\rangle\subseteq U$ and $B$ is almost disjoint from every element of $\mathcal{A}$. Then $B\in\mathcal{H}^{-}(\mathcal{A})$ and $\langle B\rangle\subseteq U$, so we’re done. Otherwise, let $\{V_{n}:n<\omega\}\subseteq\mathcal{A}$ be such that for all $n$, $\langle B_{n}\rangle\subseteq V_{n}\cap U$ for some $B_{n}\in E^{[\infty]}$. By Lemma 4.12, there exists some $B\in E^{[\infty]}$ such that:
$\langle B\rangle\subseteq\left\langle\bigcup_{n<\omega}\langle B_{n}\rangle\right\rangle\subseteq U,$
and $B$ is compatible with $B_{n}$ for all $n$. This implies that $B$ is compatible with $V_{n}$ for all $n$, so $B\in\mathcal{H}(\mathcal{A})$ and $\langle B\rangle\subseteq U$, as desired. ∎
4.3. Selectivity properties revisited
We now return to studying the combinatorial properties of the semicoideals $\mathcal{H}^{-}(\mathcal{A})$ and $\mathcal{H}(\mathcal{A})$. In Mathias’ process of proving that there are no analytic mad families of $\mathbb{N}$$^{\infty}$, he proved that if $\mathcal{A}$ is an almost disjoint family in $\mathbb{N}$$^{\infty}$, then $\mathcal{H}^{-}(\mathcal{A})$ is a selective coideal. In [24], Smythe proved that if $\mathcal{A}$ is a mad family of subspaces, then $\mathcal{H}(\mathcal{A})$ is a (p)-semicoideal. We prove a few extensions of this result.
Proposition 4.15.
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces. Then $\mathcal{H}^{-}(\mathcal{A})$ is a (p)-semicoideal.
Proof.
Let $(B_{n})_{n<\omega}$ be a $\leq$-decreasing sequence in $\mathcal{H}^{-}(\mathcal{A})$. If there exists some $C$ which weakly diagonalises $(B_{n})_{n<\omega}$ and $C$ is almost disjoint from every element of $\mathcal{A}$, then $C\in\mathcal{H}^{-}(\mathcal{A})$ and we are done, so assume otherwise. Let $C_{0}$ be a weak diagonalisation of $(B_{n})_{n<\omega}$. By our assumption, $C_{0}$ is compatible with some $V_{0}\in\mathcal{A}$, so by taking a further block subspace of $C_{0}$ if necessary, we may assume that $\langle C_{0}\rangle\subseteq V_{0}$. Now assume, for the induction hypothesis, that we have defined $C_{k},V_{k}$ for $k<m$ such that: